### v1

# Language Models use Lookbacks to Track Beliefs

Abstract
How do language models (LMs) represent characters’ beliefs, especially when those beliefs may differ from reality? This question lies at the heart of understanding the Theory of Mind (ToM) capabilities of LMs. We analyze Llama-3-70B-Instruct’s ability to reason about characters’ beliefs using causal mediation and abstraction. We construct a dataset that consists of simple stories where two characters each separately change the state of two objects, potentially unaware of each other’s actions. Our investigation uncovered a pervasive algorithmic pattern that we call a lookback mechanism, which enables the LM to recall important information when it becomes necessary. The LM binds each character-object-state triple together by co-locating reference information about them, represented as their Ordering IDs (OIs) in low rank subspaces of the state token’s residual stream. When asked about a character’s beliefs regarding the state of an object, the binding lookback retrieves the corresponding state OI and then an answer lookback retrieves the state token. When we introduce text specifying that one character is (not) visible to the other, we find that the LM first generates a visibility ID encoding the relation between the observing and the observed character OIs. In a visibility lookback, this ID is used to retrieve information about the observed character and update the observing character’s beliefs. Our work provides insights into the LM’s belief tracking mechanisms, taking a step toward reverse-engineering ToM reasoning in LMs.
1 Introduction
The ability to infer the mental states of others—known as Theory of Mind (ToM)—is an essential aspect of social and collective intelligence (Premack & Woodruff, 1978; Riedl et al., 2021). Recent studies have established that language models (LMs) can solve some tasks requiring ToM reasoning (Street et al., 2024; Strachan et al., 2024a; Kosinski, 2024), while others have highlighted shortcomings (Sclar et al., 2025; Shapira et al., 2024; Kim et al., 2023a, inter alia). Nonetheless, most existing work relies on behavioral evaluations, which do not shed light on the internal mechanisms by which LMs encode and manipulate representations of mental states to solve (or fail to solve) such tasks (Hu et al., 2025; Gweon et al., 2023).
In this work, we investigate how LMs represent and update characters’ beliefs, which is a fundamental element of ToM (Dennett, 1981; Wimmer & Perner, 1983). For instance, the Sally-Anne test (Baron-Cohen et al., 1985), a canonical test of ToM in humans, evaluates this ability by asking individuals to track Sally’s belief, which diverges from reality due to missing information, and Anne’s belief, which updates based on new observations.
We construct CausalToM, a dataset of simple stories involving two characters, each interacting with an object to change its state, with the possibility of observing one another. We then analyze the internal mechanisms that enable Llama-3-70B-Instruct (Grattafiori et al., 2024) to reason about and answer questions regarding the characters’ beliefs about the state of each object (for a sample story, see Section 3 and for the full prompt refer to Appendix A).
Figure 1: The lookback mechanism is used to perform conditional reasoning. The source token contains information that is copied into two instances via attention to create a pointer and an address. Alongside the address in the residual stream is a payload of information. When necessary, the model retrieves the payload by dereferencing the pointer. Solid lines represent movement of information, while the dotted line indicates the attention “looking back” from pointer to address.
We discover a pervasive computation that performs multiple subtasks, which we refer to as the lookback mechanism. This mechanism enables the model to recall important information only when it becomes necessary. In a lookback, two copies of a single piece of information are transferred to two distinct tokens. This allows attention heads at the latter token to look back at the earlier one when needed and retrieve vital information stored there, rather than transferring that information directly (see Fig. 1).
We identify three key lookback mechanisms that collectively perform belief tracking: 1) Binding lookback (Fig. 2(a)(i)): First the LM assigns ordering IDs (OIs) (Dai et al., 2024) that encode whether a character, object, or state token appears first or second. Then, the character and object OIs are copied to low-rank subspaces of the corresponding state token and the final token residual stream. Later, when the LM needs to answer a question about a character’s beliefs, it uses this information to retrieve the answer state OI. 2) Answer lookback (Fig. 2(a)(ii)): Uses the answer state OI from the binding lookback to retrieve the answer state token value. 3) Visibility lookback (Fig. 6): When an explicit visibility condition between characters is mentioned, the model employs additional reference information called the visibility ID to retrieve information about the observed character, augmenting the observing character’s awareness.
Overall, this work not only advances our understanding of the internal computations in LMs that enable ToM capability but also uncovers a pervasive mechanism that serves as the foundation for executing complex logical reasoning with conditionals. All code and data supporting this study are available at https://belief.baulab.info.
2 The Lookback Mechanism
((a))
Carla and Bob are working in a busy restaurant. To complete an order, Carla grabs an opaque cup and fills it with tea. Then Bob grabs another opaque bottle and fills it with water.
Question: What does Carla believe the cup contains?
Answer: tea
Bob and Carla are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee.
Question: What does Carla believe the cup contains?
Answer: coffee
Intervention 1: Answer Pointer (Refer to caption),  Causal Model Output: beer Intervention 2: Answer Payload (Refer to caption),  Causal Model Output: tea
Refer to captionCounterfactualOriginal
((b))
Figure 2: Belief Tracking in Language Models: We task the LM with tracking the beliefs of two characters that manipulate the states of two objects. We hypothesize that the LM solves this task by implementing a causal model with two lookback mechanisms (2(a)). To support our hypothesis, we conduct a causal analysis where we measure whether interventions on a high-level causal model produce the same output as equivalent interventions on the LM. For instance in 2(b), we show results for an experiment distinguishing the pointer and payload in the answer lookback.
(a) Belief Tracking with No Visibility between Characters: Our hypothesized causal model for this kind of story has two lookbacks that operate on ordering IDs (OIs) that encode whether a token appears first or second. In the binding lookback (i), the LM first represents the two events in the story by binding together each character-object-state triple in the residual stream of the state token. When questioned about a particular character and object, the LM looks back to the corresponding triple and retrieves an OI to that state token. Notice that in this lookback, that payload is later used as a pointer, i.e., what a C programmer would call a double pointer. In the answer lookback (ii), the LM dereferences the pointer to the answer token to generate the correct answer. Color indicates the information content, while shape indicates the role of that information in lookback (see Fig. 1), e.g., the state OI is a payload (Refer to caption) in the binding lookback and a pointer/address (Refer to caption) in the answer lookback.
(b) Answer Lookback Pointer and Payload: To test our hypothesized causal model, we run the LM on pairs of slightly different stories, and then intervene by patching a specific representation state from the counterfactual run to the original run, observing any change in the output. The causal model predicts that if we alter the “answer payload Refer to caption” of the original to instead take the value of the counterfactual answer payload, the output should change from coffee to tea; the gray curve in the line plot shows this does occur with $p\approx 1.0$ when patching states at the “:” token beyond layer 56, providing evidence that the answer payload resides in those states. On the other hand the causal model predicts that taking the counterfactual “answer pointer Refer to caption” would change the original run output from coffee to beer—a new output that matches neither the original nor the counterfactual!—and we do see this surprising effect, again with $p\approx 1.0$, when patching layers between 34 and 52, providing strong evidence that the answer pointer is encoded at those layers. Collected over $N=80$ samples, these measurements suggest the Answer Lookback occurs between layers 52 and 56. Furthermore the representations of the causal variables are small: the interventions can be localized even further to subspaces of dimension 33 (payload) or 20 (pointer), tiny portions of the 8192-dimensional state space.
Our investigation of belief tracking uncovers a recurring pattern of computation that we call the lookback mechanism. [^1] Although this mechanism may resemble induction heads (Elhage et al., 2021; Olsson et al., 2022), they differ fundamentally. In induction heads, information from a previous token occurrence is passed only to the subsequent token through, without being duplicated to its next occurrence. In contrast, the lookback mechanism copies the same information not only to the location where the vital information resides but also to the target location that needs to retrieve that information. Here we give a brief overview of this mechanism; subsequent sections provide detailed experiments and analyses. In lookback, source information is copied (via attention) into an address copy in the residual stream of a recalled token and a pointer copy in the residual stream of a lookback token that occurs later in the text. The LM places the address alongside a payload of the recalled token’s residual stream that can be brought forward to the lookback token if necessary. Fig. 1 schematically describes a generic lookback.
That is, the LM can use attention to dereference the pointer and retrieve the payload present in the residual stream of the recalled token (that might contain aggregated information from previous tokens), bringing it to the residual stream of the lookback token. Specifically, the pointer at the lookback token forms an attention query vector, while the address at the recalled token forms a key vector. Because the pointer and the address are copies of the same source information, they would have a high dot-product, hence a QK-circuit (Elhage et al., 2021) is established forming a bridge from the lookback token to the recalled token. The LM uses this bridge to move the payload that contains information needed to complete the subtask through the OV-circuit.
To develop an intuition for why an LM would learn to implement lookback mechanisms to solve reasoning tasks such as our belief tracking task, consider that during training LMs process text in sequence with no foreknowledge of what might come next. Then, it would be useful to mark addresses alongside payloads that might be useful for downstream tasks. In our setting, the LM constructs a representation of the story without any knowledge of what questions it may be asked about, so the LM concentrates pieces of information in the residual stream of certain tokens which later become payloads and addresses. When the question text is reached, pointers are constructed that reference this crucial story information and dereference it as the answer to the question.
3 Preliminaries
Dataset    Existing datasets for evaluating ToM capabilities of LMs are designed for behavioral testing and lack counterfactual pairs needed for causal analysis (Kim & Sundar, 2012). To address this, we constructed CausalToM, a structured dataset of simple stories, where each story involves two characters, each interacting with a distinct object causing the object to take a unique state. For example: “Character1 and Character2 are working in a busy restaurant. To complete an order, Character1 grabs an opaque Object1 and fills it with State1. Then Character2 grabs another opaque Object2 and fills it with State2.” We then ask the LM to reason about one of the characters’ beliefs regarding the state of an object: “What does Character1 believe Object2 contains?” We analyze the LM’s ability to track characters’ beliefs in two distinct settings. (1) No Visibility, where both characters are unaware of each other’s actions, and (2) Explicit Visibility where explicit information about whether a character can/cannot observe the other’s actions is provided, e.g., “Bob can observe Carla’s actions. Carla cannot observe Bob’s actions.” We also provide general task instructions (e.g., answer unknown when a character is unaware); refer to Appendix A & B for the full prompt and additional dataset details. Our experiments analyze the Llama-3-70B-Instruct model in half-precision, using NNsight (Fiotto-Kaufman et al., 2025). The model demonstrates a high behavioral performance on both the no-visibility and explicit-visibility settings, achieving accuracy of 95.7% and 99% respectively. For all subsequent experiments, we filter out samples that the model fails to answer correctly.
Causal Mediation Analysis
Our goal is to develop a mechanistic understanding of how Llama-3-70B-Instruct reasons about characters’ beliefs and answers related questions (Saphra & Wiegreffe, 2024). A key method for conducting causal analysis is interchange interventions (Vig et al., 2020; Geiger et al., 2020; Finlayson et al., 2021), in which the LM is run on paired examples: an original input $\mathbf{o}$ and a counterfactual input $\mathbf{c}$ and certain internal activations in the LM run on the original are replaced with those computed from the counterfactual.
Figure 3: Tracing information flow of crucial input tokens using causal mediation analysis.
Drawing inspiration from existing literature (Vig et al., 2020; Meng et al., 2022; Wang et al., 2023), we begin our analysis by performing interchange interventions with counterfactuals that are identical to the original except for key input tokens. We trace the causal path from these key tokens to the final output. This is a type of Causal Mediation Analysis (Mueller et al., 2024). Specifically, we construct a counterfactual dataset where $\mathbf{o}$ contains a question about the belief of a character not mentioned in the story, while $\mathbf{c}$ is identical except that the story includes the queried character. The expected outcome of this intervention is a change in the final output of $\mathbf{o}$ from unknown to a state token, such as beer. We conduct similar interchange interventions for object and state tokens (refer to Appendix C for more details).
Figure 3 presents the aggregated results of this experiment for the key input tokens Character1, Object1, and State1. The cells are color-coded to indicate the interchange intervention accuracy (IIA; Geiger et al., 2022). Even at this coarse level of analysis, several significant insights emerge: 1) Information from the correct state token (beer) flows directly from its residual stream to that of the final token in later layers, consistent with prior findings  (Lieberum et al., 2023; Prakash et al., 2024); 2) Information associated with the query character and the query object is retrieved from their earlier occurrences and passed to the final token before being replaced by the correct state token.
Desiderata Based Patching via Causal Abstraction
The causal mediation experiments provide a coarse-grained analysis of how information flows from an input token to the output but do not identify what that information is. A fact about transformers is that the input to the first layer contains input tokens and the output from the final layer contains the output token, but what is the information content of representations along the active causal path in between the input and output?
To answer this question, we turn to Causal Abstraction (Geiger et al., 2021; 2024). We align the variables of a high-level causal model with the LM’s internal activations and verify the alignment by conducting targeted interchange interventions for each variable. Specifically, we perform aligned interchange interventions at both levels: interventions that target high-level causal variables and interventions that modify low-level features of the LM’s hidden activations. If the LM produces the same output as the high-level causal model under these aligned interventions, it provides evidence supporting the hypothesized causal model. The effect of these interventions is quantified using IIA, which measures the proportion of instances where intervened high-level causal model and low-level LM have the same output (refer to Appendix D for more details about the causal abstraction framework and Appendix E for the belief tracking causal model).
In addition to performing interchange interventions on entire residual stream vectors in LMs, we also intervene on specific subspaces to further localize causal variables. To identify the subspace encoding a particular variable, we employ the Desiderata-based Component Masking (De Cao et al., 2020; Davies et al., 2023; Prakash et al., 2024) technique, which learns a sparse binary mask over the internal activation space by maximizing the logit of the causal model output token. Specifically, we train a mask to select the singular vectors of the activation space that encode a high-level variable (see Appendix F for details).
4 Belief Tracking via Ordering IDs and Lookback Mechanisms
The LM solves the no visibility setting of the belief tracking task using three key mechanisms: Ordering ID assignment, binding lookback, and answer lookback. Figure 2(a) illustrates the hypothesized high-level causal model implemented by the LM, which we evaluate in the following subsections. The LM first assigns ordering IDs (OIs; [Uncaptioned image], [Uncaptioned image], [Uncaptioned image]) to each character, object, and state in the story that encode their order of appearance (e.g., the second character Carla is assigned [Uncaptioned image]). These OIs are used in two lookback mechanisms. (i) Binding lookback: Address copies of each character OI ([Uncaptioned image]) and object OI ([Uncaptioned image]) are placed alongside their corresponding state OI payload ([Uncaptioned image]) in the residual stream of each state token, binding together each character-object-state triple. When the model is asked about the belief of a specific character about a specific object, it moves pointer copies of the corresponding OIs ([Uncaptioned image], [Uncaptioned image]) to the final token’s residual stream. These pointers are dereferenced, bringing the correct state OI into the final token residual stream. (ii) Answer lookback: An address copy of the state OI ([Uncaptioned image]) is alongside the state token payload ([Uncaptioned image]) in the residual stream of the correct state token, while a pointer copy ([Uncaptioned image]) is moved to the final token residual stream via the binding lookback. The pointer is dereferenced, bringing the answer state token payload into the final token residual stream, which is predicted as the final output.
Refer to Appendix E for pseudocode defining the causal model for the belief tracking task. In Appendices L and K, we show that parts of our analysis generalizes to the Llama-3.1-405B-Instruct model and the BigToM dataset (Gandhi et al., 2024).
4.1 Ordering ID Assignment
The LM assigns an Ordering ID (OI; Dai et al. 2024) to the character, object, and state tokens. These OIs, encoded in a low-rank subspace of the internal activation, serve as a reference that indicates whether an entity is the first or second of its type independent of its token value. For example, in Fig. 2(a), Bob is assigned the first character OI, while Carla receives the second. In the subsequent subsections and Appendices G & H, we validate the presence of OIs through multiple experiments, where intervening on tokens with identical token values but different OIs alters the model’s internal computation, leading to systematic changes in the final output predicted by our high-level causal model. The LM then uses these OIs as building blocks, feeding them into lookback mechanisms to track and retrieve beliefs.
4.2 Uncovering the Binding Lookback Mechanism
Carla and Bob are working in a busy restaurant. To complete an order, Carla grabs an opaque cup and fills it with coffee. Then Bob grabs another opaque bottle and fills it with beer.
Question: What does Carla believe the cup contains?
Answer: coffee
Bob and Carla are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee.
Question: What does Carla believe the cup contains?
Answer: coffee
Intervention: Binding Payloads and Addresses (Refer to caption, Refer to caption, Refer to caption) Causal Model Output: beer
Refer to captionCounterfactualOriginal
Figure 4: Binding lookback payload and address: We intervened on both the high-level causal model and the LM running on the original story, modifying their variables and internal activations respectively, to match those from a counterfactual scenario. In the causal model intervention, we update the addresses (character and object OIs; Refer to caption and Refer to caption) and the payloads (state OIs; Refer to caption). This causes the binding lookback mechanism to attend to and retrieve the state OI corresponding to the alternate state token, which is then dereferenced by answer lookback to yield the alternate state token (e.g., beer instead of coffee). In the LM interchange intervention, modifying the residual stream at the state token results in identical outputs between layers 33 and 38. This confirms our hypothesis that both the address and payload information are represented in the residual stream of state tokens.
The Binding Lookback is the first operation applied to these OIs. The character and object OIs, serving as the source information, are each copied twice. One copy, referred to as the address, is placed in the residual stream of the state token (recalled token), alongside the state OI as the payload to transfer. The other copy, referred to as the pointer, is moved in the residual stream of the final token (lookback token). These pointer and address copies are then used to form the QK-circuit at the lookback token, which dereferences the state OI payload, transferring it from the state token to the final token. See Fig.2(a) (i) for a schematic of this lookback and see Fig.1 for the general mechanism.
Localizing the Address and Payload    In our first experiment, we localize the address copies of the character and object OIs and the state OI payload to the residual stream of the state token (recalled token, Fig. 2(a)). We sampled a counterfactual dataset where each example consists of an original input $\mathbf{o}$ with an answer that isn’t unknown and a counterfactual input $\mathbf{c}$ where the character, object, and state tokens are identical, except the ordering of the two sentences is swapped while the question remains unchanged, as illustrated in Fig. 4. The expected outcome predicted by our high-level causal model under intervention is the other state token from the original example, e.g., beer, because reversing the address and payload values without changing the pointer flips the output. In LM, the QK-circuit, formed using the pointer at the lookback token, attends to the other state token and retrieves its state OI as the payload.
We perform an interchange intervention experiment layer-by-layer, where we replace the residual stream vectors at the first state token in the original run with that of the second state token in the counterfactual run and vice versa for the other state token. It is important to note that if the intervention targets state token values instead of their OIs, it should not produce the expected output. (This happens in the earlier layers.)
As shown in Fig. 4, the strongest alignment occurs between layers $33$ and $38$, supporting our hypothesis that the state token’s residual stream contains both the address information (character and object OIs) and the payload information (state OI). These components are subsequently used to form the appropriate QK and OV circuits of binding lookback.
Localizing the Source Information    As shown in Fig. 2(a), the source information is copied as both the address and the pointer at different token positions. To localize the source information, we conduct intervention experiments with a dataset where the counterfactual example, c, swaps the order of the characters and objects as well as replaces the state tokens with entirely new ones while keeping the question the same as in o.
With this dataset, an interchange intervention on the high-level causal model that targets the source information will have downstream effects on both the address and the pointer, so no change in output occurs. However, if we additionally freeze the payloads and addresses, the causal model outputs the other state token, e.g., beer in Fig. 5, due to the mismatch between address and pointer.
In the LM, we interchange the residual streams of the character and object tokens while keeping the residual stream of the state token fixed. When the output of the intervened LM aligns with that of the intervened causal model, it indicates that the QK-circuit at the final token is attending to the alternate state token. As shown in Fig. 5, the second experiment reveals alignment between layers 20 and 34. This suggests that source information—specifically, the character and object OIs—is represented in their respective token residual streams within this layer range.
Carla and Bob are working in a busy restaurant. To complete an order, Carla grabs an opaque cup and fills it with tea. Then Bob grabs another opaque bottle and fills it with water.
Question: What does Carla believe the cup contains?
Answer: tea
Bob and Carla are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee.
Question: What does Carla believe the cup contains?
Answer: coffee
Intervention: Binding Source (Refer to caption, Refer to caption) Frozen: Binding and Answer Addresses+Payloads (Refer to caption, Refer to caption, Refer to caption; Refer to caption, Refer to caption) Causal Model Output: beer
Refer to captionCounterfactualOriginal
Figure 5: Source Information of Binding lookback: We run the causal model and the LM on an original story, then update their variables and activations, respectively, to the values they would take on for a counterfactual story with swapped characters and objects and new states. The interchange intervention on the high-level causal model swaps the sources of the binding lookback (character and object OIs; Refer to caption, Refer to caption) while freezing the addresses and payloads of the binding lookback (character, object, and state OIs; Refer to caption, Refer to caption, Refer to caption) and the answer lookback (state OI and token; Refer to caption, Refer to caption). By altering the sources, but freezing the addresses and payloads, only the pointer is changed so the binding lookback retrieves the other state OI which is dereferenced by the answer lookback to the other state token (e.g., beer instead of coffee). We perform the same interchange intervention on the LM and measure the agreement with the intervened causal model. Our results localize the source to the character and object token residual streams between layers 20 and 34.
We provide more experimental results in Appendix G where we show in Fig. 12 that freezing the residual stream of the state token is necessary. In sum, these results not only provide evidence for the presence of source information but also establish its transfer to the recalled and lookback tokens as addresses and pointers, respectively.
Localizing the Pointer Information    The pointer copies of the character and object OI are first formed at the character and object tokens in the question before being moved again to the final token for dereferencing (see Appendix H for experiments and more details).
4.3 Uncovering the Answer Lookback Mechanism
The LM answers the question using the Answer Lookback. The state OI of the correct answer serves as the source information, which is copied into two instances. One instance, the address copy of the state OI, is in the residual stream of the state token (recalled token) with the state token itself as the payload. The other instance, the pointer copy of the state OI, is transferred to the residual stream of the final token (lookback token) as the payload of the binding lookback. This pointer is then dereferenced, bringing the state token as the payload into the residual stream of the final token, which is predicted as the final output. See Fig. 2(a) (ii) for the answer lookback and Fig. 1 for the general mechanism.
Localizing the Pointer Information    We first localize the pointer of the answer lookback, which is the payload of the binding lookback. To do this, we conduct an interchange intervention experiment where the residual vectors at the final token position in the original run are replaced with those from the counterfactual run, one layer at a time. The counterfactual inputs have swapped objects and characters and randomly sampled states. If the answer pointer is targeted for intervention in the high-level causal model, the output is the other state in the original input, e.g., beer. As shown in Fig. 2(b), alignment begins at layer $34$, indicating that this layer contains pointer information, in low-rank subspace, which remains causally relevant until layer $52$.
Localizing the Payload    To determine where the model uses the state OI pointer to retrieve the state token, we use the same interchange intervention experiment. However, if the answer payload is targeted for intervention in the high-level causal model, the output is the correct state token from the counterfactual example, e.g., tea, rather than the state token from the original example, as illustrated in Fig. 2(b). The alignment occurs after layer $56$, indicating that the model retrieves the correct state token (payload) into the final token’s residual stream by $56$, where it is subsequently used to generate the final output.
5 Impact of Visibility Conditions on Belief Tracking Mechanism
In the previous section, we demonstrated how the LM uses ordering IDs and two lookback mechanisms to track the beliefs of characters that cannot observe each other. Now, we explore how the LM updates the beliefs of characters when provided with additional information that one of the characters (observing) can observe the actions of others (observed). We hypothesize that the LM employs another lookback mechanism, which we refer to as the Visibility Lookback, to incorporate information about the observed character.
As illustrated in Fig. 6, we hypothesize that the LM first generates a Visibility ID at the residual stream of the visibility sentence, serving as the source information. The address copy of the visibility ID remains in the residual stream of the visibility sentence, while its pointer copy gets transferred to the residual streams of the subsequent tokens, which are the lookback tokens. Then LM forms a QK-circuit at the lookback tokens and dereferences the visibility ID pointer to bring forward the payload.
Figure 6: Visbility Lookback When one character (the observing character) can see another (the observed character), the LM assigns a visibility ID (Refer to caption) to the visibility sentence (where this relation is defined). An address copy of this visibility ID remains in the visibility sentence’s residual stream. A pointer copy of the visibility ID is transferred to the subsequent tokens’ residual stream (lookback tokens). During processing, the model dereferences this pointer through a QK-circuit, bringing forward the payload ( Refer to caption ). Based on initial evidence, this payload contains the observed character’s OI(Refer to caption). Refer to Appendix I for more details. This mechanism allows the model to incorporate the observed character’s knowledge into the observing character’s belief state, enabling more complex belief reasoning.
Although we were unable to determine the exact semantics of the payload in this lookback, we speculate that it encodes the observed character’s OI. We propose the existence of another lookback, where the story sentence associated with the observed character serves as the source, and its payload encodes information about the observed character. That payload information, encoding observed character’s OI, is then retrieved by the lookback tokens of the Visibility lookback, which contributes to the queried character’s enhanced awareness (see Appendix I for more details).
5.1 Uncovering the Visibility Lookback Mechanism
Localizing the Source Information    To localize the source information, we conduct an interchange intervention experiment where the counterfactual is a different story with altered visibility information. In the original example, the first character cannot observe the second character’s actions, whereas in the counterfactual example, the first character can observe them (Fig. 7). The causal model outcome of this intervention is a change in the final output of the original run from “unknown” to the state token associated with the queried object. The interchange intervention is executed on visibility sentence tokens. As shown in Fig. 7 ( line), alignment occurs between layers $10$ and $23$, indicating that the visibility ID remains encoded in the visibility sentence until layer $23$, after which it is duplicated into address and pointer copies on visibility sentence and subsequent tokens respectively.
Localizing the Payload    To localize the payload information, we use the same counterfactual dataset. However, instead of intervening on the source or recalled tokens, we intervene on the lookback tokens, specifically the question and answer tokens. As in the previous experiment, we replace the residual vectors of these tokens in the original run with those from the counterfactual run. As shown in Fig. 7 ( line), alignment occurs only after layer $31$, indicating that the information enhancing the queried character’s awareness is present in the lookback tokens only after this layer.
Carla and Bob are working in a busy restaurant. To complete an order, Carla grabs an opaque cup and fills it with tea. Then Bob grabs another opaque bottle and fills it with water. Bob cannot observe Carla’s actions. Carla can observe Bob’s actions.
Question: What does Carla believe the bottle contains?
Answer: water
Karen and Max are working in a busy restaurant. To complete an order, Karen grabs an opaque flute and fills it with soda. Then Max grabs another opaque jar and fills it with coffee. Max cannot observe Karen’s actions. Karen cannot observe Max’s actions.
Question: What does Karen believe the jar contains?
Answer: unknown
Intervention 1: Source (Refer to caption), Causal Model Output: coffee Intervention 2: Payload ( Refer to caption ), Causal Model Output: coffee Intervention 3: Address & Pointer (Refer to caption), Causal Model Output: coffee
Refer to captionCounterfactualOriginal
Figure 7: Visibility Lookback: We conduct three interchange intervention experiments to support the Visibility Lookback hypothesis: (1) Source Alignment: We align the source information (Refer to caption) by intervening on the visibility sentence—replacing it with its representation from a counterfactual run where the visibility sentence causes the queried character to become aware of the queried object’s contents. We observe that source information aligns between layers $10$ and $23$, after which it splits into separate address and pointer components. (2) Payload Alignment: To align the payload ( Refer to caption ), we intervene on all subsequent tokens and observe alignment only after layer $31$. (3) Address and Pointer Alignment: When intervening on both the address and pointer information (Refer to caption), we observe alignment across a broader range of layers, particularly between layers $24$ and $31$, because of the enhanced alignment between the address and pointer copies at the recalled and lookback tokens.
Localizing the Address and Pointer    The previous two experiments suggest the presence of a lookback mechanism, as there is no signal indicating that the source or payload has been formed between layers 24 and 31. We hypothesize that this lack of signal is due to a mismatch between the address and pointer information at the recalled and lookback tokens. Specifically, when intervening only on the recalled token after layer $25$, the pointer is not updated, whereas intervening only on the lookback tokens leaves the address unaltered, leading to the mismatch. To test this hypothesis, we conduct another intervention using the same counterfactual dataset, but this time, we intervene on the residual vectors of both the recalled and lookback tokens, i.e., the visibility sentence, as well as the question and answer tokens. As shown in Fig. 7 ( line), alignment occurs after layer $10$ and remains stable, supporting our hypothesis. This intervention replaces both the address and pointer copies of the visibility IDs, enabling the LM to form a QK-circuit and retrieve the payload.
6 Related Work
Theory of mind in LMs    A large body of work has focused on benchmarking different aspects of ToM through various tasks that attempt to assess LMs’ performance such as Le et al. (2019); Xu et al. (2024); Shapira et al. (2023); Jin et al. (2024); Wu et al. (2023); Kim et al. (2023b); Chan et al. (2024); Strachan et al. (2024b) and many more. In addition, there are various methods tailored to improve ToM ability in LMs through prompting (e.g., Sclar et al., 2023; Zhou et al., 2023; Wilf et al., 2024; Moghaddam & Honey, 2023; Hou et al., 2024).
Entity tracking in LMs    Entity tracking and variable binding are crucial abilities for LMs to exhibit not only coherent ToM capabilities, but also neurosymbolic reasoning. Many existing works have attempted to decipher this ability in LMs (Li et al., 2021; Davies et al., 2023; Kim & Schuster, 2023; Prakash et al., 2024; Feng & Steinhardt, 2023; Feng et al., 2024; Dai et al., 2024). Our work builds on their empirical insights and extends the current understanding of how LMs bind various entities defined in context.
Mechanistic interpretability of theory of mind    Only a few empirical studies explored the underlying mechanisms of ToM of LM (Zhu et al., 2024; Bortoletto et al., 2024) (Herrmann & Levinstein, 2024, is a notable theoretical paper). Those studies focus on probing techniques (Belinkov, 2022; Alain, 2016) to identify internal representations of beliefs and used steering techniques (Li et al., 2024; Rimsky et al., 2023) to improve LM performance by manipulating their activations. However, the mechanism by which LMs solve those tasks remains a black box, limiting our ability to understand, predict, and control LMs’ behaviors.
7 Conclusion
Through a series of desiderata-based patching experiments, we have mapped the mechanisms underlying the processing of partial knowledge and false beliefs in a set of simple stories. We are surprised by the pervasive appearance of a single recurring computational pattern: the lookback, which resembles a pointer dereference inside a transformer. The LMs use a combination of several lookbacks to reason about nontrivial visibility and belief states. Our improved understanding of these fundamental computations gives us optimism that it may be possible to fully reveal the algorithms underlying not only Theory of Mind, but also other forms of reasoning in LMs.
8 Acknowledgement
This research was supported in part by Open Philanthropy (N.P., N.S., A.S.S., D.B., A.G., Y.B.), the NSF National Deep Inference Fabric award #2408455 (D.B.), the Israel Council for Higher Education (N.S.), the Zuckerman STEM Leadership Program (T.R.S.), the Israel Science Foundation (grant No. 448/20; Y.B.), an Azrieli Foundation Early Career Faculty Fellowship (Y.B.), a Google Academic Gift (Y.B.), and a Google Gemma Academic Award (D.B.). This research was partly funded by the European Union (ERC, Control-LM, 101165402). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Full prompt
No Visibility Instruction: 1. Track the belief of each character as described in the story. 2. A character’s belief is formed only when they perform an action themselves or can observe the action taking place. 3. A character does not have any beliefs about the container and its contents which they cannot observe. 4. To answer the question, predict only what is inside the queried container, strictly based on the belief of the character, mentioned in the question. 5. If the queried character has no belief about the container in question, then predict ‘unknown’. 6. Do not predict container or character as the final output.
Story: Bob and Carla are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee.
Question: What does Bob believe the bottle contains?
Answer:
Explicit Visibility Instruction: 1. Track the belief of each character as described in the story. 2. A character’s belief is formed only when they perform an action themselves or can observe the action taking place. 3. A character does not have any beliefs about the container and its contents which they cannot observe. 4. To answer the question, predict only what is inside the queried container, strictly based on the belief of the character, mentioned in the question. 5. If the queried character has no belief about the container in question, then predict ‘unknown’. 6. Do not predict container or character as the final output.
Story: Bob and Carla are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee. Bob can observe Carla’s actions. Carla cannot observe Bob’s actions.
Question: What does Bob believe the cup contains?
Answer:
Appendix B The CausalToM Dataset
In total, there are 4 templates (one without and 3 with explicit visibility statements). Each template allows 4 different types of questions (CharacterX asked about ObjectY). We used lists of 103 characters, 21 objects, and 23 states. In our interchange intervention experiments, we randomly sample 80 pairs of original and counterfactual stories.
Appendix C Causal Mediation Analysis
Bob and Carla are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee.
Question: What does Bob believe the bottle contains?
Answer: beer
David and Carla are working in a busy restaurant. To complete an order, David grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee.
Question: What does Bob believe the bottle contains?
Answer: unknown
Causal Model Output: beer
CounterfactualOriginal
Figure 8: Causal Mediation Analysis: The original example produces the output unknown because Bob is not mentioned in the story, leaving the model without any information about his beliefs. However, when the residual stream vectors corresponding to Bob from the counterfactual run are patched into the original run, the model acquires the necessary information about that character and consequently updates its output to beer.
In addition to the experiment shown in Fig.8, we conduct similar experiments for the object and state tokens by replacing them in the story with random tokens, which alters the original example’s final output. However, patching the residual stream vectors of these tokens from the counterfactual run restores the relevant information, enabling the model to predict the causal model output. The results of these experiments are collectively presented in Fig.3, with separate heatmaps shown in Fig. 91011.
Figure 9: Information flow of character input tokens using causal mediation analysis.
Figure 10: Information flow of object input tokens using causal mediation analysis.
Figure 11: Information flow of state input tokens using causal mediation analysis.
Appendix D Desiderate Based Patching Via Causal Abstraction
Causal Models and Interventions    A deterministic causal model $\mathcal{M}$ has variables that take on values. Each variable has a mechanism that determines the value of the variable based on the values of parent variables. Variables without parents, denoted $\mathbf{X}$, can be thought of as inputs that determine the setting of all other variables, denoted $\mathcal{M}(\mathbf{x})$. A hard intervention $A\leftarrow a$ overrides the mechanisms of variable $A$, fixing it to a constant value $a$.
Interchange Interventions    We perform interchange interventions (Vig et al., 2020; Geiger et al., 2020) where a variable (or set of features) $A$ is fixed to be the value it would take on if the LM were processing counterfactual input $\mathbf{c}$. We write $A\leftarrow\mathsf{Get}(\mathcal{M}(\mathbf{c}),A)$ where $\mathsf{Get}(\mathcal{M}(\mathbf{c}),A)$ is the value of variable $A$ when $\mathcal{M}$ processes input $\mathbf{c}$. In experiments, we will feed a original input $\mathbf{o}$ to a model under an interchange intervention $\mathcal{M}_{A\leftarrow\mathsf{Get}(\mathcal{M}(\mathbf{c}),A))}(\mathbf{o})$.
Featurizing Hidden Vectors    The dimensions of hidden vectors are not an ideal unit of analysis (Smolensky, 1986), and so it is typical to featurize a hidden vector using some invertible function, e.g., an orthogonal matrix, to project a hidden vector into a new variable space with more interpretable dimensions called “features”(Mueller et al., 2024). A feature intervention $\mathbf{F}_{\mathbf{h}}\leftarrow\mathbf{f}$ edits the mechanism of a hidden vector $\mathbf{h}$ to fix the value of features $\mathbf{F}_{\mathbf{h}}$ to $\mathbf{f}$.
Alignment    The LM is a low-level causal model $\mathcal{L}$ where variables are dimensions of hidden vectors and the hypothesis about LM structure is a high-level causal model $\mathcal{H}$. An alignment $\Pi$ assigns each high-level variable $A$ to features of a hidden vector $\mathbf{F}^{A}_{\mathbf{h}}$, e.g., orthogonal directions in the activation space of $\mathbf{h}$. To evaluate an alignment, we perform intervention experiments to evaluate whether high-level interventions on the variables in $\mathcal{H}$ have the same effect as interventions on the aligned features in $\mathcal{L}$.
Causal Abstraction    We use interchange interventions to reveal whether the hypothesized causal model $\mathcal{H}$ is an abstraction of an LM $\mathcal{L}$. To simplify, assume both models share an input and output space. The high-level model $\mathcal{H}$ is an abstraction of the low-level model $\mathcal{L}$ under a given alignment when each high-level interchange intervention and the aligned low-level intervention result in the same output. For a high-level intervention on $A$ aligned with low-level features $\mathbf{F}^{A}_{\mathbf{h}}$ with a counterfactual input $\mathbf{c}$ and original input $\mathbf{b}$, we write
$\mathsf{GetOutput}(\mathcal{L}_{\mathbf{F}_{\mathbf{h}}^{A}\leftarrow\mathsf{Get}(\mathcal{L}(\mathbf{c}),\mathbf{F}^{A}_{\mathbf{h}}))}(\mathbf{o}))=\mathsf{GetOutput}(\mathcal{H}_{A\leftarrow\mathsf{Get}(\mathcal{H}(\mathbf{c}),A))}(\mathbf{o}))$ (1)
If the low-level interchange intervention on the LM produces the same output as the aligned high-level intervention on the algorithm, this is a piece of evidence in favor of the hypothesis. This extends naturally to multi-variable interventions (Geiger et al., 2024).
Graded Faithfulness Metric    We construct counterfactual datasets for each causal variable where an example consists of a base prompt and a counterfactual prompt . The counterfactual label is the expected output of the algorithm after the high-level interchange intervention, i.e., the right-side of Equation 1. The interchange intervention accuracy is the proportion of examples for which Equation 1 holds, i.e., the degree to which $\mathcal{H}$ faithfully abstracts $\mathcal{L}$.
Appendix E Pseudocode for the Belief Tracking High-Level Causal Model
Algorithm 1 Belief Representation
1:procedure BeliefTracking($c_{1},o_{1},s_{1},c_{2},o_{2},s_{2},q_{c},q_{o}$)
2:  Ordering ID assignment
3:  $c_{1}^{OI},o_{1}^{OI},s_{1}^{OI}\leftarrow\text{AssignOIs}([c_{1},o_{1},s_{1}],1)$
4:  $c_{2}^{OI},o_{2}^{OI},s_{2}^{OI}\leftarrow\text{AssignOIs}([c_{2},o_{2},s_{2}],2)$
5:
6:  Binding lookback mechanism
7:  $\text{binding\_address}_{1}\leftarrow(\text{copy}(c_{1}^{OI}),\text{copy}(o_{1}^{OI}))$
8:  $\text{binding\_address}_{2}\leftarrow(\text{copy}(c_{2}^{OI}),\text{copy}(o_{2}^{OI}))$
9:
10:  $q_{c}^{OI}\leftarrow\text{copy}(\{c_{1}:c_{1}^{OI},c_{2}:c_{2}^{OI}\}[q_{c}])$
11:  $q_{o}^{OI}\leftarrow\text{copy}(\{o_{1}:o_{1}^{OI},o_{2}:o_{2}^{OI}\}[q_{o}])$
12:  $\text{binding\_pointer}\leftarrow(q_{c}^{OI},q_{o}^{OI})$
13:
14:  if $\text{binding\_address}_{1}=\text{binding\_pointer}$ then
15:   $\text{binding\_payload}\leftarrow\text{copy}(s_{1}^{OI})$
16:  else if $\text{binding\_address}_{2}=\text{binding\_pointer}$ then
17:   $\text{binding\_payload}\leftarrow\text{copy}(s_{2}^{OI})$
18:  end if
19:
20:  Answer lookback mechanism
21:  $\text{answer\_pointer}\leftarrow\text{binding\_payload}$
22:  $\text{answer1\_address}\leftarrow s_{1}^{OI}$
23:  $\text{answer2\_address}\leftarrow s_{2}^{OI}$
24:  if $\text{answer1\_address}=\text{answer\_pointer}$ then
25:   $\text{answer\_payload}\leftarrow s_{1}$
26:  else if $\text{answer2\_address}=\text{answer\_pointer}$ then
27:   $\text{answer\_payload}\leftarrow s_{2}$
28:  end if
29:  return answer_payload
30:end procedure
Algorithm 2 High-level causal model for the no visibility
Appendix F Desiderata-based Component Masking
While interchange interventions on residual vectors reveal where a causal variable might be encoded in the LM’s internal activations, they do not localize the variable to specific subspaces. To address this, we apply the Desiderata-based Component Masking technique (De Cao et al., 2020; Davies et al., 2023; Prakash et al., 2024), which learns a sparse binary mask over the singular vectors of the LM’s internal activations. First, we cache the internal activations from $500$ samples at the token positions where residual-level interchange interventions align with the expected output. Next, we apply Singular Value Decomposition to compute the singular vectors, which are then used to construct a projection matrix. Rather than replacing the entire residual vector with that from the counterfactual run, we perform subspace-level interchange interventions using the following equations:
$\mathsf{W_{proj}=V*V^{T}}$ (2)
$\mathsf{h_{org}\leftarrow W_{proj}h_{counterfactual}+(I-W_{proj})h_{original}}$ (3)
Here, $V$ is a matrix containing stacked singular vectors, while $h_{counterfactual}$ and $h_{original}$ represent the residual stream vectors from the counterfactual and original runs, respectively. The core idea is to first remove the existing information from the subspace defined by the projection matrix and then insert the counterfactual information into that same subspace using the same projection matrix. However, in DCM, instead of utilizing the entire internal activation space, we learn a binary mask over the matrix containing stacked singular vectors to identify the desired subspace. Specifically, before computing the projection matrix, we use the following equations to select the relevant singular vectors:
$\mathsf{V\leftarrow V*mask}$ (4)
We train the mask on $80$ examples of the same counterfactual dataset and use another $80$ as the validation set. We use the following objective function, which maximizes the logit of the causal model output token:
$\mathcal{L}=-\mathsf{logit_{expected\_output}}+\lambda\sum\mathsf{W}$ (5)
Where $\lambda$ is a hyperparameter used to control the rank of the subspace and $\mathsf{W}$ is the parameter of the learnable mask. We trained it for one epoch with ADAM optimizer, a batch size of $4$ and a learning rate of $0.01$.
Appendix G Aligning Character and Object OIs
As mentioned in section 4, the source information, consisting of character and object OI, is duplicated to form the address and pointer of the binding lookback. Here, we describe another experiment to verify that the source information is copied to both the address and the pointer. More specifically, we conduct the same interchange intervention experiment as described in Fig. 5, but without freezing the residual vectors at the state tokens. Based our hypothesis, this intervention will not be able to change the state of the original run, since the intervention at the source information will affect both address and pointer, hence making the model form the original QK-circuit.
Carla and Bob are working in a busy restaurant. To complete an order, Carla grabs an opaque cup and fills it with coffee. Then Bob grabs another opaque bottle and fills it with beer.
Question: What does Carla believe the cup contains?
Answer: coffee
Bob and Carla are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee.
Question: What does Carla believe the cup contains?
Answer: coffee
Intervention: Binding Source (Refer to caption, Refer to caption) Causal Model Output: beer
Refer to captionCounterfactualOriginal
Figure 12: Source Information of Binding lookback: In this interchange intervention experiment, the source information—i.e., the character and object OIDs (Refer to caption, Refer to caption)—is modified, while the address and payload (Refer to caption, Refer to caption, Refer to caption) are recomputed based on the modified source. Since both the address and pointer information are derived from the altered source, the binding lookback ultimately retrieves the same original state token as the payload. As a result, we do not observe high intervention accuracy.
In section 4, we identified the source of the information but did not fully determine the locations of each character and object OI. To address this, we now localize the character and object OIs separately to gain a clearer understanding of the layers at which they appear in the residual streams of their respective tokens, as shown in Fig.13 and Fig.14.
Bob and Carla are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee.
Question: What does Carla believe the cup contains?
Answer: coffee
Carla and Bob are working in a busy restaurant. To complete an order, Carla grabs an opaque cup and fills it with tea. Then Bob grabs another opaque bottle and fills it with water.
Question: What does Carla believe the bottle contains?
Answer: unknown
Intervention: Character OI (Refer to caption) Frozen: Object OI, Binding and Answer Addresses+Payloads (Refer to caption, Refer to caption, Refer to caption; Refer to caption, Refer to caption) Causal Model Output: water
Refer to captionCounterfactualOriginal
Figure 13: Character OI: This interchange intervention experiment swaps the character OI (Refer to caption), while freezing the object OI as well as binding lookback address and payload (Refer to caption, Refer to caption, Refer to caption). Swapping the character OIs in the story tokens changes the queried character OI to the other one. Hence, the final output changes from unknown to water.
Bob and Carla are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee.
Question: What does Carla believe the cup contains?
Answer: coffee
Carla and Bob are working in a busy restaurant. To complete an order, Carla grabs an opaque cup and fills it with tea. Then Bob grabs another opaque bottle and fills it with water.
Question: What does Carla believe the bottle contains?
Answer: unknown
Intervention: Binding Source (Refer to caption, Refer to caption) Frozen: Binding and Answer Addresses+Payloads, Queried Character OID (Refer to caption, Refer to caption, Refer to caption; Refer to caption, Refer to caption) Causal Model Output: tea
Refer to captionCounterfactualOriginal
Figure 14: Object OI: This interchange intervention experiment swaps both the character and object OIs (Refer to caption, Refer to caption), while freezing the address and payload of binding lookback (Refer to caption, Refer to caption, Refer to caption) as well as queried character OI (Refer to caption). Swapping both character and object OIs in the story tokens ensures that the queried object gets the other OI. Hence, the final output changes from unknown to tea.
Appendix H Aligning Query Character and Object OIs
In section 4.2, we localized the pointer information of binding lookback. However, we found that this information is transferred to the lookback token (last token) through two intermediate tokens: the queried character and the queried object. In this section, we separately localize the OIs of the queried character and queried object, as shown in Fig. 15 and Fig. 16.
Bob and Carla are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee.
Question: What does Carla believe the cup contains?
Answer: coffee
Carla and Bob are working in a busy restaurant. To complete an order, Carla grabs an opaque cup and fills it with tea. Then Bob grabs another opaque bottle and fills it with water.
Question: What does Carla believe the bottle contains?
Answer: unknown
Intervention: Binding Pointer (Refer to caption) Causal Model Output: water
Refer to captionCounterfactualOriginal
Figure 15: Query Character OI: This interchange intervention experiment alters the OI of the queried character (Refer to caption) to the other one. Hence, the final output changes from unknown to water.
Bob and Carla are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee.
Question: What does Carla believe the cup contains?
Answer: coffee
Carla and Bob are working in a busy restaurant. To complete an order, Carla grabs an opaque cup and fills it with tea. Then Bob grabs another opaque bottle and fills it with water.
Question: What does Bob believe the cup contains?
Answer: unknown
Intervention: Binding Pointer (Refer to caption) Causal Model Output: water
Refer to captionCounterfactualOriginal
Figure 16: Query Object OI: This interchange intervention experiment alters the OI of the queried object (Refer to caption) to the other one. Hence, the final output changes from unknown to water.
Appendix I Speculated Payload in Visibility Lookback
As mentioned in section 5, the payload of the Visibility lookback remains undetermined. In this section, we attempt to disambiguate its semantics using the Attention Knockout technique introduced in (Geva et al., 2023), which helps reveal the flow of crucial information. We apply this technique to understand which previous tokens are vital for the formation of the payload information. Specifically, we ”knock out” all attention heads at all layers of the second visibility sentence, preventing them from attending to one or more of the previous sentences. Then, we allow the attention heads to attend to the knocked-out sentence one layer at a time.
If the LM is fetching vital information from the knocked-out sentence, the interchange intervention accuracy (IIA) post-knockout will decrease. Therefore, a decrease in IIA will indicate which attention heads, at which layers, are bringing in the vital information from the knocked-out sentence. If, however, the model is not fetching any critical information from the knocked-out sentence, then knocking it out should not affect the IIA.
Figure 17: At the second visibility sentence, attention heads are restricted to retrieve information from one of three prior contexts: (1) both the second story sentence and the first visibility sentence ( line), (2) only the first visibility sentence ( line), or (3) only the second story sentence ( line).
To determine if any vital information is influencing the formation of the Visibility lookback payload, we perform three knockout experiments: 1) Knockout attention heads from the second visibility sentence to both the first visibility sentence and the second story sentence (which contains information about the observed character), 2) Knockout attention heads from the second visibility sentence to only the first visibility sentence, and 3) Knockout attention heads from the second visibility sentence to the second story sentence. In each experiment, we measure the effect of the knockout using IIA.
Fig.17 shows the experimental results. Knocking out any of the previous sentences affects the model’s ability to produce the correct output. The decrease in IIA in the early layers can be explained by the restriction on the movement of character OIs. Specifically, the second visibility sentence mentions the first and second characters, whose character OIs must be fetched before the model can perform any further operations. Therefore, we believe the decrease in IIA until layer 15, when the character OIs are formed (based on the results from Section G), can be attributed to the model being restricted from fetching the character OIs. However, the persistently low IIA even after this layer—especially when both the second and first visibility sentences are involved—indicates that some vital information is being fetched by the second visibility sentence, which is essential for forming the coherent Visibility lookback payload. Thus, we speculate that the Visibility payload encodes information about the observed character, specifically their character OI, which is later used to fetch the correct state OI.
Appendix J Correlation Analysis of Causal Subspaces and Attention Heads
This section identifies the attention heads that align with the causal subspaces discovered in the previous sections. Specifically, first we focus on attention heads whose query projections are aligned with the subspaces—characterized by the relevant singular vectors—that contain the correct answer state OI. To quantify this alignment between attention heads and causal subspaces, we use the following computation.
Let $Q\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ denote the query projection weight matrix for a given layer:
We normalize $Q$ column-wise:
$\tilde{Q}_{:,j}=\frac{Q_{:,j}}{\|Q_{:,j}\|}\quad\text{for each column }j$ (6)
Let $S\in\mathbb{R}^{d_{\text{model}}\times k}$ represent the matrix of $k$ singular vectors (i.e., the causal subspace basis). We project the normalized query weights onto this subspace:
$Q_{\text{sv}}=\tilde{Q}\cdot S$ (7)
We then reshape the resulting projection into per-head components. Assuming $Q_{\text{sv}}\in\mathbb{R}^{d_{\text{model}}\times k}$, and each attention head has dimensionality $d_{h}$, we write:
$Q_{\text{head}}^{(i)}=Q_{\text{sv}}^{(i)}\in\mathbb{R}^{d_{h}\times k}\quad\text{for }i=1,\dots,n_{\text{heads}}$ (8)
Finally, we compute the norm of each attention head’s projection:
$\text{head\_norm}_{i}=\left\|Q_{\text{head}}^{(i)}\right\|_{F}\quad\text{for }i=1,\dots,n_{\text{heads}}$ (9)
We compute the head_normhead\_norm for each attention head in every layer, which quantifies how strongly a given head reads from the causal subspace present in the residual stream. The results are presented in Fig. 18, and they align with our previous findings: attention heads in the later layers form the QK-circuit by using pointer and address information to retrieve the payload during the Answer lookback.
We perform a similar analysis to check which attention heads’ value projection matrix align with the causal subspace that encodes the payload of the Answer lookback. Results are shown in Fig. 19, indicating that attention heads at later layers primarily align with causal subspace containing the answer token.
Figure 18: Alignment between the Answer lookback pointer causal subspace and the query projection matrix in Llama-3-70B-Instruct.
Figure 19: Alignment between the Answer lookback payload causal subspace and the value projection matrix in Llama-3-70B-Instruct.
Appendix K Belief Tracking Mechanism in BigToM Benchmark
This section presents preliminary evidence that the mechanisms outlined in Sections 4 and 5 generalize to other benchmark datasets. Specifically, we demonstrate that Llama-3-70B-Instruct answers the belief questions (true belief and false belief) in the BigToM dataset Gandhi et al. (2024) in a manner similar to that observed for CausalToM: by first converting token values to their corresponding OIs and then performing logical operations on them using lookbacks. However, as noted in Section 3, BigToM—like other benchmarks—lacks the coherent structure necessary for causal analysis. As a result, we were unable to replicate all experiments conducted on CausalToM. Thus, the results reported here provide only preliminary evidence of a similar underlying mechanism.
Noor is working as a barista at a busy coffee shop. Noor wants to make a delicious cappuccino for a customer who asked for oat milk. Noor grabs a milk pitcher and fills it with oat milk. A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task. Noor sees her coworker swapping the milk.
Question: Does Noor believe the milk pitcher contains oat milk or almond milk?
Answer: almond milk
Tariq is a fisherman in a coastal village in Indonesia. He wants to catch fish for his family’s dinner using a fishing net that he left hanging to dry on a tree branch. Tariq sees tat the net appears to be dry and ready to use. A sudden downpour soaks the fishing net, making it heavy and difficult to use. Tariq notices the rain soaking the fishing net.
Question: Does Noor believe the fishing net is dry or soaked?
Answer: unknown
Causal Model Output: soaked
Refer to captionCounterfactualOriginal
Figure 20: Query Character OI in BigToM: This interchange intervention experiment inserts the first character’s OI into the residual stream at the queried character token (Refer to caption), resulting in the movement of pointer information to the last token that aligns with the address information of binding lookback mechanism. Consequently, the model is able to form the appropriate QK-circuit from the last token to predict the correct state answer token(s) as the final output, instead of unknown.
To justify the presence of OIs, we conduct an interchange intervention experiment, similar to the one described in Section H, aiming to localize the character OI at the character token in the question sentence. We construct an original sample by replacing its question sentence with that of a counterfactual sample, selected directly from the unaltered BigToM dataset. Consequently, when processing the original sample, the model has no information about the queried character and, as a result, produces unknown as the final output. However, if we replace the residual vector at the queried character token in the original sample with the corresponding vector from the counterfactual sample (which contains the character OI), the model’s output changes from unknown to the state token(s) associated with the queried object. This is because inserting the character OI at the queried token provides the correct pointer information, aligning with the address information at the correct state token(s), thereby enabling the model to form the appropriate QK-circuit and retrieve the state’s OI. As shown in Fig. 20, we observe a high IIA between layers 9289-28—similar to the pattern seen in CausalToM—suggesting that the queried character token encodes the character OI in its residual vector within these layers.
Next, we investigate the Answer lookback mechanism in BigToM, focusing specifically on localizing the pointer and payload information at the final token position. To localize the pointer information, which encodes the correct state OI, we construct original and counterfactual samples by selecting two completely different examples from the BigToM dataset, each with different ordered states as the correct answer. For example, as illustrated in Fig.21, the counterfactual sample designates the first state as the answer, thrilling plot, whereas the original sample designates the second state, almond milk. We perform an intervention by swapping the residual vector at the last token position from the counterfactual sample into the original run. The causal model outcome of this intervention is that the model will output the alternative state token from the original sample, oat milk. As shown in Fig.21, this alignment occurs between layers 33 and 51, similar to the layer range observed for the pointer information in the Answer lookback of CausalToM.
Yael is in a small bookstore browsing through the science fiction section. Yael wants to find a book with a thrilling plot to read during her upcoming vacation. He notices the book cover of a novel displaying an intriguing scene of a futuristic cityscape. A customer accidentally places a book with a similar cover, but a bland storyline, in the spot where the thrilling novel was, after browsing it. Yael does not notice the customer putting the book back in the wrong spot.
Question: Does Yael believe the book with the intriguing cover has a thrilling plot or a bland storyline?
Answer: thrilling plot
Noor is working as a barista at a busy coffee shop. Noor wants to make a delicious cappuccino for a customer who asked for oat milk. Noor grabs a milk pitcher and fills it with oat milk. A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task. Noor sees her coworker swapping the milk.
Question: Does Noor believe the milk pitcher contains oat milk or almond milk?
Answer: almond milk
Causal Model Output: oat milk
Refer to captionCounterfactualOriginal
Figure 21: Answer Lookback Pointer in BigToM: This interchange intervention experiment modifies the pointer information (Refer to caption) of the Answer lookback, thereby altering the subsequent QK-circuit to attend to the other state (e.g., oat milk) instead of the original one (e.g., almond milk). As a result, the model retrieves the token value corresponding to the other state to answer the question.
Yael is in a small bookstore browsing through the science fiction section. Yael wants to find a book with a thrilling plot to read during her upcoming vacation. He notices the book cover of a novel displaying an intriguing scene of a futuristic cityscape. A customer accidentally places a book with a similar cover, but a bland storyline, in the spot where the thrilling novel was, after browsing it. Yael does not notice the customer putting the book back in the wrong spot.
Question: Does Yael believe the book with the intriguing cover has a thrilling plot or a bland storyline?
Answer: thrilling plot
Noor is working as a barista at a busy coffee shop. Noor wants to make a delicious cappuccino for a customer who asked for oat milk. Noor grabs a milk pitcher and fills it with oat milk. A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task. Noor sees her coworker swapping the milk.
Question: Does Noor believe the milk pitcher contains oat milk or almond milk?
Answer: almond milk
Causal Model Output: thrilling plot
Refer to captionCounterfactualOriginal
Figure 22: Answer Lookback Payload in BigToM: This interchange intervention experiment directly modifies the payload information (Refer to caption) of the Answer lookback, which is fetched from the corresponding state tokens and predicted as the next token(s). Thus, replacing its value in the original run, e.g. almond milk, with that from the counterfactual run, e.g. thrilling plot, causes the model’s next predicted tokens to correspond to the correct answer of the counterfactual sample.
Further, to localize the payload of the Answer lookback in BigToM, we perform an interchange intervention experiment using the same original and counterfactual samples as mentioned in the previous experiment, but with a different expected output—namely, the correct state from the counterfactual sample instead of the other state from the original sample. As shown in Fig. 22, alignment emerges after layer 59, consistent with the layer range observed for the Answer lookback payload in CausalToM.
Noor is working as a barista at a busy coffee shop. Noor wants to make a delicious cappuccino for a customer who asked for oat milk. Noor grabs a milk pitcher and fills it with oat milk. A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task. Noor sees her coworker swapping the milk.
Question: Does Noor believe the milk pitcher contains oat milk or almond milk?
Answer: almond milk
Noor is working as a barista at a busy coffee shop. Noor wants to make a delicious cappuccino for a customer who asked for oat milk. Noor grabs a milk pitcher and fills it with oat milk. A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task. Noor does not see her coworker swapping the milk.
Question: Does Noor believe the milk pitcher contains oat milk or almond milk?
Answer: oak milk
Expected Output: almond milk
Refer to captionCounterfactualOriginal
Figure 23: Visibility Lookback in BigToM: We perform three interchange interventions to establish the presence of the Visibility ID, which serves as both address and pointer information. When intervening at the source (Refer to caption)—i.e., the visibility sentence—both the address and pointer are updated, resulting in alignment across layers. Intervening only at the subsequent question tokens leads to alignment only at later layers, after the model has already fetched the payload ( Refer to caption ). However, intervening at both the visibility and question sentences results in alignment across all layers, as the address and pointer remain consistent throughout.
Finally, we investigate the impact of the visibility condition on the underlying mechanism and find that, similar to CausalToM, the model uses the Visibility lookback to enhance the observing character’s awareness based on the observed character’s actions. To localize the effect of the visibility condition, we perform an interchange intervention in which the original and counterfactual samples differ in belief type—that is, if the original sample involves a false belief, the counterfactual involves a true belief, and vice versa. The expected output of this experiment is the other (incorrect) state of the original sample. Following the methodology in Section 5, we conduct three types of interventions: (1) only at the visibility condition sentence, (2) only at the subsequent question sentence, and (3) at both the visibility condition and the question sentence. As shown in Fig. 23, intervening only at the visibility sentence results in alignment at early layers, up to layer $17$, while intervening only at the subsequent question sentence leads to alignment after layer $26$. Intervening on both the visibility and question sentences results in alignment across all layers. These results align with those found in the CausalToM setting shown in the Fig. 7.
Previous experiments suggest that the underlying mechanisms responsible for answering belief questions in BigToM are similar to those in CausalToM. However, we observed that the subspaces encoding various types of information are not shared between the two settings. For example, although the pointer information in the Answer lookback encodes the correct state’s OI in both cases, the specific subspaces that represent this information at the final token position differ significantly. We leave a deeper investigation of this phenomenon—shared semantics across distinct subspaces in different distributions—for future work.
Appendix L Generalization of Belief Tracking Mechanism on CausalToM to Llama-3.1-405B-Instruct
This section presents all the interchange intervention experiments described in the main text, conducted using the same set of counterfactual examples on Llama-3.1-405B-Instruct, using NDIF  Fiotto-Kaufman et al. (2025). Each experiment was performed on 80 samples. Due to computational constraints, subspace interchange intervention experiments were not conducted. The results indicate that Llama-3.1-405B-Instruct employs the same underlying mechanism as Llama-3-70B-Instruct to reason about belief and answer related questions. This suggests that the identified belief-tracking mechanism generalizes to other models capable of reliably performing the task.
Carla and Bob are working in a busy restaurant. To complete an order, Carla grabs an opaque cup and fills it with coffee. Then Bob grabs another opaque bottle and fills it with beer.
Question: What does Carla believe the cup contains?
Answer: coffee
Bob and Carla are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee.
Question: What does Carla believe the cup contains?
Answer: coffee
Causal Model Output: beer
Refer to captionCounterfactualOriginal
Figure 24: Payload and address of Binding lookback
Carla and Bob are working in a busy restaurant. To complete an order, Carla grabs an opaque cup and fills it with tea. Then Bob grabs another opaque bottle and fills it with water.
Question: What does Carla believe the cup contains?
Answer: tea
Bob and Carla are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee.
Question: What does Carla believe the cup contains?
Answer: coffee
Causal Model Output: beer
Refer to captionCounterfactualOriginal
Figure 25: Source Information of Binding lookback
Carla and Bob are working in a busy restaurant. To complete an order, Carla grabs an opaque cup and fills it with coffee. Then Bob grabs another opaque bottle and fills it with beer.
Question: What does Carla believe the cup contains?
Answer: coffee
Bob and Carla are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee.
Question: What does Carla believe the cup contains?
Answer: coffee
Causal Model Output: beer
Refer to captionCounterfactualOriginal
Figure 26: Source Information of Binding lookback without freezing address and payload
Bob and Carla are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee.
Question: What does Carla believe the cup contains?
Answer: coffee
Carla and Bob are working in a busy restaurant. To complete an order, Carla grabs an opaque cup and fills it with tea. Then Bob grabs another opaque bottle and fills it with water.
Question: What does Carla believe the bottle contains?
Answer: unknown
Causal Model Output: water
Refer to captionCounterfactualOriginal
Figure 27: Character OI
Bob and Carla are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee.
Question: What does Carla believe the cup contains?
Answer: coffee
Carla and Bob are working in a busy restaurant. To complete an order, Carla grabs an opaque cup and fills it with tea. Then Bob grabs another opaque bottle and fills it with water.
Question: What does Carla believe the bottle contains?
Answer: unknown
Causal Model Output: tea
Refer to captionCounterfactualOriginal
Figure 28: Object OI
Bob and Carla are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee.
Question: What does Carla believe the cup contains?
Answer: coffee
Carla and Bob are working in a busy restaurant. To complete an order, Carla grabs an opaque cup and fills it with tea. Then Bob grabs another opaque bottle and fills it with water.
Question: What does Bob believe the cup contains?
Answer: unknown
Causal Model Output: water
Refer to captionCounterfactualOriginal
Figure 29: Query Object OI
Bob and Carla are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee.
Question: What does Carla believe the cup contains?
Answer: coffee
Carla and Bob are working in a busy restaurant. To complete an order, Carla grabs an opaque cup and fills it with tea. Then Bob grabs another opaque bottle and fills it with water.
Question: What does Carla believe the bottle contains?
Answer: unknown
Causal Model Output: water
Refer to captionCounterfactualOriginal
Figure 30: Query Character OI
Carla and Bob are working in a busy restaurant. To complete an order, Carla grabs an opaque cup and fills it with tea. Then Bob grabs another opaque bottle and fills it with water.
Question: What does Carla believe the cup contains?
Answer: tea
Bob and Carla are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee.
Question: What does Carla believe the cup contains?
Answer: coffee
Intervention: Answer Pointer (Refer to caption),  Causal Model Output: beer Intervention: Answer Payload (Refer to caption),  Causal Model Output: tea
Refer to captionCounterfactualOriginal
Figure 31: Answer Lookback Pointer and Payload
Carla and Bob are working in a busy restaurant. To complete an order, Carla grabs an opaque cup and fills it with tea. Then Bob grabs another opaque bottle and fills it with water. Bob cannot observe Carla’s actions. Carla can observe Bob’s actions.
Question: What does Carla believe the bottle contains?
Answer: water
Bob and Carla are working in a busy restaurant. To complete an order, Carla grabs an opaque cup and fills it with beer. Then Bob grabs another opaque bottle and fills it with coffee. Bob cannot observe Carla’s actions. Carla cannot observe Bob’s actions.
Question: What does Carla believe the bottle contains?
Answer: unknown
Causal Model Output: coffee
Refer to captionCounterfactualOriginal
Figure 32: Visibility Lookback